# Noise Matters: Optimizing Matching Noise for Diffusion Classifiers

### v1

@@ Appendix A Background of Diffusion Classifiers (DC)
The Implementations of DC. In practice, DC implementations often utilize pretrained text-to-image diffusion models such as Stable Diffusion Rombach et al. (2022); Ramesh et al. (2022). The timestep weights can be even or varied, such as based on the signal-to-noise ratio (SNR). Another typical implementation involves computing a re-weighted reconstruction error across diffusion time-steps, selecting the class whose conditioning prompt results in minimal reconstruction error at an early time-step Yue et al. (2024). More efficient implementations reduce computational overhead by applying variance reduction strategies, including shared-noise sampling and candidate class pruning. Alternatively, few-shot DC learn class-specific concepts on top of pretrained diffusion models to enhance the classification of nuanced attributes, improving robustness against spurious visual correlations Yue et al. (2024).
The advantages of DC. Compared with vision-language model (VLM) classifiers such as CLIP, DC exhibits several significant advantages. Firstly, DC demonstrates superior multimodal compositional reasoning and attribute binding capabilities, outperforming VLM classifiers on tasks requiring nuanced understanding of visual features and their textual descriptions Clark & Jaini (2023); Li et al. (2023). Secondly, DC inherently possesses robustness to texture-shape biases and spurious correlations, a common issue in discriminative few-shot training with VLMs, by utilizing a principled generative modeling approach that isolates semantic attributes effectively through diffusion time-steps Yue et al. (2024). Lastly, the generative modeling nature of DC allows seamless adaptation to few-shot and zero-shot learning settings without extensive retraining, making them versatile and efficient for various downstream tasks Clark & Jaini (2023); Yue et al. (2024).
@@ Appendix B Theory of 2-D Fourier Transform
The Fourier transform decomposes an image into its constituent frequency components, effectively distinguishing low-frequency (smooth or slowly varying) and high-frequency (sharp edges and fine details) content Yin et al. (2019); Rao et al. (2021). In our frequency matching validation (c.f., Sec. 4.7), leveraging this decomposition provides valuable insights into how diffusion processes interact with different frequency components, informing analysis and manipulation strategies at various frequency scales Yu et al. (2022). After the above frequency decomposing technique, for each image or noise latent, we can represent its frequency by its high-frequency signals ratio. Specifically, we set a cutoff threshold (e.g., 0.3 was used for Figure 7(b)) to divide all the signals into two parts: high-frequency and low-frequency. The ratio of high-frequency signals to low-frequency signals is used to represent its frequency.
@@ D.1 Additional Few-shot Learning Comparisons
Settings. To further evaluate the effectiveness of our NoOp, we conducted additional few-shot (shot=1,2,4,8) learning experiments. Specifically, we compared our NoOp with two baselines, i.e., ensemble (sample five different noises for one image) zero-shot DC Li et al. (2023); Clark & Jaini (2023) and prompt learning DC Yue et al. (2024); Zhou et al. (2022b) (c.f., Section 3.1) across eights datasets: CIFAR-10 Krizhevsky et al. (2009), CIFAR-100 Krizhevsky et al. (2009), Flowers102 Nilsback & Zisserman (2008), DTD Cimpoi et al. (2014), OxfordPets Parkhi et al. (2012), EuroSAT Helber et al. (2019), STL-10 Coates et al. (2011) and FGVCAircraft Maji et al. (2013). We used Stable Diffusion-v2.0 Ramesh et al. (2022) as the pretrained diffusion model. Results are averaged on three random seeds.
@@
D.2 Effectiveness of NoOp across timesteps
Settings. In the main paper, we fixed the timestep at 500 to focus on the noise optimization. To evaluate whether our NoOp is consistently effective on other timesteps. We conducted a timestep ablation study. We tested $t=300,400,500,600,700$ with Stable Diffusion-v2.0 Ramesh et al. (2022) as the pretrained diffusion model on 4-shot CIFAR-10 Krizhevsky et al. (2009). Results are averaged on three random seeds.
Results. As shown in Figure 8(a), our NoOp can gain consistent improvements across different timesteps. This indicates that NoOp is robust and effective across timesteps.
D.3 Timesteps Ensemble
Settings. In zero-shot DC Li et al. (2023); Clark & Jaini (2023), there is an ensemble across timesteps to reduce the noise instability and increase the performance. To evaluate whether our NoOp can also benefit from the timestep ensemble after optimizing the noises. We test the ensemble with 5 different timesteps (i.e., $t=300,400,500,600,700$) after using NoOp. The pretrained diffusion model is Stable Diffusion-v2.0 Ramesh et al. (2022) and dataset is 4-shot CIFAR-10 Krizhevsky et al. (2009). We also showed the results of zero-shot w/o ensemble, zero-shot w/ensemble, and NoOp w/o ensemble for comparison. Results are averaged on three random seeds.
Results. As shown in Figure 7(b), we have two observations: 1) Compared with the one timestep (i.e., $t=500$) zero-shot DC, the timesteps ensemble (5 different timesteps) can improve around $23\%$ top-1 accuracy. This indicates the ensemble can effectively improve the performance with linearly increased computation overhead as a price. 2) Compared with the one timestep (i.e., $t=500$) NoOp, the timesteps ensemble can also improve around $12\%$ top-1 accuracy and outperformed the zero-shot ensemble $5.4\%$. This demonstrates that the ensemble strategy can also be applied for our NoOp if there are enough computational resources.
D.4 Meta-Network Ablation
@@
D.5 The Stability of NoOp
Settings. To evaluate the stability of NoOp, we randomly sampled three noises (three different seeds) from the Gaussian distribution as the initial noises. Then we conducted the zero-shot classification with them on CIFAR-10 Krizhevsky et al. (2009) test set, respectively. Then we used NoOp to optimize them on the 8-shot training set and test. The pretrained diffusion model we used is Stable Diffusion-v2.0 Ramesh et al. (2022). We reported the top-1 accuracies before and after training. All other settings are the same as Sec. 4.1.
@@ Appendix E Implementation Details
Details of NoOp. We used the Discrete Euler as the timestep scheduler, max length padding for the text tokenizer across all the experiments. The initial noise $\epsilon$ is randomly sampled from the standard Gaussian distribution, while the Meta-Network is a U-Net Ronneberger et al. (2015) architecture with 3 up-sampling layers and 3 down-sampling layers. Each sampling layer consists of 2-D convolutional layers with ReLu activation Agarap (2018) and BatchNorm Ioffe & Szegedy (2015). During training, we used fixed learning rates (w/o warm up and scheduler). The training batch size is 32.
Other Reproduction Details. For prompt optimization DC (c.f., Sec. 3.1) implementation in Sec. 4.2 and Sec. D.1, we followed the implementations of TiF Learner Yue et al. (2024) and textual inversion Gal et al. (2022) respectively.

### v2

@@ Appendix A Background of Diffusion Classifiers (DC)
The recent surge of visual generation benefits from Diffusion models [ho2020denoising, song2020score], and high-quality images and videos are generated by sampling from Gaussian noise. Meanwhile, the downstream tasks include editing [zhang2023adding, lu2024robust], composing [lu2023tf], and erasing [lu2024mace] are also frequently researched.
@@
The Implementations of DC. In practice, DC implementations often utilize pretrained text-to-image diffusion models such as Stable Diffusion [rombach2022high, ramesh2022hierarchical]. The timestep weights can be even or varied, such as based on the signal-to-noise ratio (SNR). Another typical implementation involves computing a re-weighted reconstruction error across diffusion time-steps, selecting the class whose conditioning prompt results in minimal reconstruction error at an early time-step [yue2024few]. More efficient implementations reduce computational overhead by applying variance reduction strategies, including shared-noise sampling and candidate class pruning. Alternatively, few-shot DC learn class-specific concepts on top of pretrained diffusion models to enhance the classification of nuanced attributes, improving robustness against spurious visual correlations [yue2024few].
The advantages of DC. Compared with vision-language model (VLM) classifiers such as CLIP, DC exhibits several significant advantages. Firstly, DC demonstrates superior multimodal compositional reasoning and attribute binding capabilities, outperforming VLM classifiers on tasks requiring nuanced understanding of visual features and their textual descriptions [clark2023text, li2023your]. Secondly, DC inherently possesses robustness to texture-shape biases and spurious correlations, a common issue in discriminative few-shot training with VLMs, by utilizing a principled generative modeling approach that isolates semantic attributes effectively through diffusion time-steps [yue2024few]. Lastly, the generative modeling nature of DC allows seamless adaptation to few-shot and zero-shot learning settings without extensive retraining, making them versatile and efficient for various downstream tasks [clark2023text, yue2024few].
@@ Appendix B Theory of 2-D Fourier Transform
The Fourier transform decomposes an image into its constituent frequency components, effectively distinguishing low-frequency (smooth or slowly varying) and high-frequency (sharp edges and fine details) content [yin2019fourier, rao2021global]. In our frequency matching validation (c.f., Sec. 4.1), leveraging this decomposition provides valuable insights into how diffusion processes interact with different frequency components, informing analysis and manipulation strategies at various frequency scales [yu2022deep]. After the above frequency decomposing technique, for each image or noise latent, we can represent its frequency by its high-frequency signals ratio. Specifically, we set a cutoff threshold (e.g., 0.3 was used for Figure 7(b)) to divide all the signals into two parts: high-frequency and low-frequency. The ratio of high-frequency signals to low-frequency signals is used to represent its frequency.
@@ D.1 Additional Few-shot Learning Comparisons
Settings. To further evaluate the effectiveness of our NoOp, we conducted additional few-shot (shot=1,2,4,8) learning experiments. Specifically, we compared our NoOp with two baselines, i.e., ensemble (sample five different noises for one image) zero-shot DC [li2023your, clark2023text] and prompt learning DC [yue2024few, zhou2022learning] (c.f., Section 3.1) across eights datasets: CIFAR-10 [krizhevsky2009learning], CIFAR-100 [krizhevsky2009learning], Flowers102 [nilsback2008automated], DTD [cimpoi2014describing], OxfordPets [parkhi2012cats], EuroSAT [helber2019eurosat], STL-10 [coates2011analysis] and FGVCAircraft [maji2013fine]. We used Stable Diffusion-v2.0 [ramesh2022hierarchical] as the pretrained diffusion model. Results are averaged on three random seeds.
@@
D.2 Meta-Network Ablation
@@
D.3 The Stability of NoOp
Settings. To evaluate the stability of NoOp, we randomly sampled three noises (three different seeds) from the Gaussian distribution as the initial noises. Then we conducted the zero-shot classification with them on CIFAR-10 [krizhevsky2009learning] test set, respectively. Then we used NoOp to optimize them on the 8-shot training set and test. The pretrained diffusion model we used is Stable Diffusion-v2.0 [ramesh2022hierarchical]. We reported the top-1 accuracies before and after training. All other settings are the same as Sec. 4.8.
@@ Appendix E Implementation Details
Details of NoOp. We used the Discrete Euler as the timestep scheduler, max length padding for the text tokenizer across all the experiments. The initial noise $\epsilon$ is randomly sampled from the standard Gaussian distribution, while the Meta-Network is a U-Net [ronneberger2015u] architecture with 3 up-sampling layers and 3 down-sampling layers. Each sampling layer consists of 2-D convolutional layers with ReLu activation [agarap2018deep] and BatchNorm [ioffe2015batch]. During training, we used fixed learning rates (w/o warm up and scheduler). The training batch size is 32.
Other Reproduction Details. For prompt optimization DC (c.f., Sec. 3.1) implementation in Sec. 4.2 and Sec. D.1, we followed the implementations of TiF Learner [yue2024few] and textual inversion [gal2022image] respectively.